A projectile is lunched with 30m/s at an angle 60 degree with the horizontal. Find the position of the projectile and the magnitude and direction of its velocity at t=2s


1
Expert's answer
2021-01-25T13:55:00-0500

An object moving in the two-dimensional parabolic path is a projectile motion. It moves in that trajectory because it is affected only by gravitational acceleration. Projectile Motion is described by two-dimensional kinematic equations.

Position

"x_t = x_o+v_0cos(\theta_o)t"

"y_t = y_o + v_osin(\theta_o)t"

Velocity

"v_x = v_ocos(\theta_o)"

"v_y = v_osin(\theta_o)-gt"

The given in the problem:

"v_0=30\frac{m}{s} \space \space \space \space \theta_0=60^o \space\space\space\space g = 9.8 \frac{m}{s^2}"

"v_x = 30*cos60^o = 15\frac{m}{s} \space\space\space\space\space v_y =15*sin60^0-9.8*2 = -6.6\frac{m}{s}"

The negative answer indicates it is moving downward.

Using Pythagoras: "v =\sqrt{v_x^2+v_y^2} = \sqrt{15^2+(-6.6)^2} = 16.4\frac{m}{s^2}"